OpenStudy (anonymous):

Sam's plasma TV has a 14 in by 18 in screen with a plastic trim border. The total area of the TV and border is 480 in2. How wide is the trim around the TV?

OpenStudy (anonymous):

Area of the plastic border is 480-(14)(18)=228. We know the length of the top/bottom and the sides of the border...same as the TV dimensions. We want 228=2(14*w)+2(18*w) where w=width of the border (this assumes rectangular border). 228=64w and w=3.56 The width of the border is about 3.56 inches

OpenStudy (anonymous):

actually I think this might be wrong..we would still have the four corners of the frame to account for. I think it should be: 228=64w + 4w^2 Solving this quadratic give w=3 inches This should be correct.